OpenStudy (anonymous):

integral x^3/(x^2-2) dx I did u sub and got (x^2-2)/2 + ln(x^2-2) + C but wolfram alpha says it's x^2/2 + ln(x^2-2) + C did I do something wrong? does u sub here don't work?

OpenStudy (zarkon):

those answers are the same

OpenStudy (anonymous):

how? I can see you could separate the fraction for the second one but where does the constant -2/2 go?

OpenStudy (zarkon):

into the other constant

OpenStudy (anonymous):

wouldn't that make the two constant different? +C and +C-1?

OpenStudy (zarkon):

\[(x^2-2)/2 +\ln(x^2-2) + C\] \[=x^2/2 +\ln(x^2-2) + C-1\] \[=x^2/2 +\ln(x^2-2) + C_2\]

OpenStudy (anonymous):

oh so because we don't know the value of original constant, we can assume any changes in the constant would not change the function?

OpenStudy (zarkon):

It is not that we don't know the original constant...the antiderivative of a function \(f\) is a Function \(F\) such that \(F'=f\) both functions above satisfy this.

OpenStudy (anonymous):

oh I think I see now thanks :)